OpenStudy (anonymous):

Solve the equation by factoring: x^4 - x = 0

jimthompson5910 (jim_thompson5910):

\[\Large x^4 - x = 0\] \[\Large x(x^3 - 1) = 0\] \[\Large x(x^3 - 1^3) = 0\] \[\Large x(x - 1)(x^2+1*x+1^2) = 0\] \[\Large x(x - 1)(x^2+x+1) = 0\] I'll let you finish. Note: I factored \(\Large x^3 - 1\) using the difference of cubes factoring rule.